# Efficient Paths for Local Counterdiabatic Driving

### v1

@@ III.2 Ansatz for Extra Controls
corresponding to the terms $H_{c}^{(1)}$ and $H_{c}^{(2)}$ in Eq. (10). Like in the previous example we then use a standard CD protocol and find the parameters $\beta^{(1)}$ and $\beta^{(2)}$ by optimizing the fidelity of the final state.
As shown in Figure  4 for the long-range Ising model, we again find a dramatic improvement of the GHZ state preparation along the optimally augmented path. We notice that this improvement is not tied in any way to the integrability (short range)/nonintegrability (long range) of the Ising model. This is not that surprising, as the short range AGP (small $\ell$) cannot distinguish integrable and nonintegrable systems Pandey et al. (2020) and extra controls generally break integrability anyway.
@@
Here we rescale the first term so that the model has a well defined classical/thermodynamic limit as $S=N/2\to\infty$. We note in passing that this Hamiltonian is extensively employed for experimental preparation of spin squeezed states Kitagawa and Ueda (1993); Law et al. (2001); Rojo (2003); Sorensen and Molmer (2001). Such spin-squeezed states allow for better scaling of measurement precision in Ramsey interferometry, surpassing the Standard Quantum Limit Wineland et al. (1992); Bollinger et al. (1996). We show the results for this in Figure 5. Interestingly one of the emergent extra control Hamiltonians in Eq. 10 is $S_{y}^{2}$, which is used in the two-axis twisting Hamiltonian for the preparation of even better spin-squeezed states.
As in both the short- and long-range spin models, we observe that we can prepare states with much better fidelity using these more efficient paths, though the improvement is less dramatic. The performance can be likely enhanced further by considering finite energy norms for the variational optimization of the action, as was done in a different classical model Gjonbalaj et al. (2022).
@@ V Conclusions & Outlook
We have proposed a systematic method for finding extra control terms for quantum annealing protocols with which to augment a Hamiltonian, so that local CD driving prepares a target ground state with exponentially better fidelity. Phrased another way, our method specifies how to find the paths through the space of possible Hamiltonian couplings along which approximate CD driving will be most efficient. We find these paths by restricting the additional terms to those which preserve the commutator structure of the original CD driving term. These terms can be effectively engineered via Floquet protocols using only the original controls. We have tested this method on 1D spin chains with short- and long-range interactions and showed that it allows for strong improvement of preparation of GHZ states.
While this work has been principally concerned with proposing this technique and testing it with simple models, there are many others to which this recipe for more efficient paths might be applied. Going beyond one-dimensional systems, exploring models with frustration, or mappings to e.g. satisfiability problems on arbitrary graphs are all areas where this recipe may yield improvement.
@@ Acknowledgments
This work was supported by NSF Grant DMR-2103658 and the AFOSR Grant FA9550-21-1-0342. The authors thank Aashish Clerk, Andrew Daley, Callum Duncan, Michael Flynn and Tatsuhiko Ikeda for helpful discussions. The authors acknowledge that the computational work reported in this paper was performed on the Shared Computing Cluster administered by Boston University’s Research Computing Services. The numerical computations were performed using QuSpin Weinberg and Bukov (2017, 2019).
@@ Appendix A Construction of AGP in Krylov space
Let us provide more detail about how the approximate AGP is constructed. In particular, we use a Krylov space construction of the AGP which is slightly different to that described in Ref. Claeys et al., 2019. The main different is that we demand that subsequent terms in the commutator expansion be orthogonal to each other. In particular, we write the AGP in terms of Krylov space operators $O_{k}$ with coefficients $\gamma_{k}$ (see also Refs. Bhattacharjee, 2023; Takahashi and del Campo, 2023)
@@
where $D=2^{N}$ is the dimension of the Hilbert space, and $H$ is the Hamiltonian. We construct the Krylov space operators $O_{k}$ according to the following algorithm:
@@
$S_{\lambda}=1+2\gamma_{1}b_{1}+\sum_{k=1}^{\ell}(\gamma_{k}(b_{2k}^{2}+b_{2k-1}^{2})+2b_{2k}b_{2k+1}\gamma_{k}\gamma_{k+1})$
@@
00 0 $=\gamma_{k}A_{k}+\gamma_{k+1}B_{k}+\gamma_{k-1}B_{k-1}+b_{1}\delta_{k,1}$
$A_{k}$ $=b_{2k}^{2}+b_{2k+1}^{2}$
@@
Appendix B Performance of YY vs. ZXZ controls in Short-Range Model
@@
In Figure 6, we show a contour plot of the final fidelity for different combinations of $\beta_{YY}$ and $\beta_{ZXZ}$ using local CD driving with $\ell=1$ and $\ell=2$ as defined in Eq. (4) for the Hamiltonian above with $N=10$. We find that for $\ell=1$ the optimal protocol is very close to one where only one of the controls is used, i.e. where either $\beta_{YY}=0$ or $\beta_{ZXZ}=0$. The situation reverses for $\ell=2$, where the optimal performance is achieved for $\beta_{YY}\approx-\beta_{ZXZ}$. Figure 3 in the main text shows the protocol performance along the corresponding optimal directions.
Figure 6: The improvement obtained by augmenting the short-range GHZ state preparation for $N=10$ by either $YY$ or $ZXZ$ controls. The top shows $\ell=1$ an the bottom shows $\ell=2$. The x and y axis show the coefficient for the $YY$ and $ZXZ$ controls respectively. From the symmetry of this plot, it is clear that neither has an advantage over the other. Starting at $\ell=2$, using both simultaneously provides a significant advantage.
Appendix C Details of the Method
In this section, we will describe in detail the step-by-step process by which we find these improved paths for local CD driving (see also Ref. Čepaitė et al., 2023). The first step is to determine the Hamiltonian for the physical system of interest, and represent it as a matrix. Furthermore, it is useful to leverage any symmetries present in the system and so construct the Hamiltonian in the relevant symmetry sector. This will make the local CD driving more efficient by reducing the number of transitions it tries to suppress. In the language of Eq. (5), this means determining in $H_{0}$ or $H_{1}$ and then forming
@@ Appendix C Details of the Method
In this paper, we are working far from the adiabatic limit, so as long as $|\dot{\lambda}|\to\infty$, the precise way that the time-dependence is encoded in $\lambda$ does not matter. However, for completeness we mention that we use a protocol which vanishes smoothly at $\lambda=0$ and $\lambda=1$:
$\lambda(t)=\sin^{2}\left(\frac{\pi}{2}\sin^{2}\left(\frac{\pi t}{2\tau}\right)\right)$
We then compute the fidelity using Eq. (8). We employ the Powell minimization algorithm Powell (1964) to choose the values of $\beta^{(1)}$ and $\beta^{(2)}$ which maximize the fidelity. Practically, we limit the values the coefficients can take to $|\beta^{(i)}|<3$. This concludes the steps to implementing our protocol. In Figure 7 we apply this procedure to the short-range model Hamiltonian of Eq. (9).
Figure 7: Improvement in the final state fidelity obtained by using the extra control ansatz of Eq. (10) to prepare a GHZ state using the short-range Hamiltonian of Eq. (9). This can be compared with Figure 3 where we use $H_{c}^{(1)}=YY$ and $H_{c}^{(2)}=ZXZ$ for the same protocol. Note that they are extremely similar because the corresponding extra controls are linear combinations of each other up to a rescaling of $ZZ$ and $X$ in the original Hamiltonian. As before, red points indicate following the “naive” original path, whereas blue indicates following an “augmented” path.
@@
Appendix D Explanation of Different Regimes of Fidelity Improvement
Figure 8: The variational optimization in terms of fitting excitation frequencies, for the short-range model with $N=6$. On top is the “naive path” of the original Hamiltonian, and on the bottom is the augmented path. Although this is just one snapshot, it is apparent from this that the extra control Hamiltonian $H_{c}=YY$ makes local CD driving more efficient by bringing multiple excitation frequencies together.
As remarked upon in the main text, in Figures 3, 4 and 7, we see two types of behavior for the fidelity of the final state prepared with extra control Hamiltonians. The first regime is a plateau where we can prepare the final state with near unit fidelity up to a certain system size, whereas the second regime has less than unit fidelity but is still exponentially improved when compared to the “naive” path.

### v2

@@ IV.1 Fidelity Gain from Augmented Hamiltonian
corresponding to the terms $H_{c}^{(1)}$ and $H_{c}^{(2)}$ in Eq. (12). Like in the previous example we then use a standard CD protocol and find the parameters $\beta^{(1)}$ and $\beta^{(2)}$ by numerically optimizing the fidelity of the final state.
In Figure  5, for the long-range Ising model, we find a dramatic improvement in the GHZ state preparation along the optimally augmented path. We notice that this improvement is not tied in any way to the integrability (short range)/nonintegrability (long range) of the Ising model. This is not that surprising, as the short range AGP (small $\ell$) cannot distinguish integrable and nonintegrable systems Pandey et al. (2020) and extra controls generally break integrability anyway.
@@
Here we rescale the first term so that the model has a well defined classical/thermodynamic limit as $S=N/2\to\infty$. We note in passing that this Hamiltonian is extensively employed for experimental preparation of spin squeezed states Kitagawa and Ueda (1993); Law, Ng, and Leung (2001); Rojo (2003); Sorensen and Molmer (2001). Such spin-squeezed states allow for better scaling of measurement precision in Ramsey interferometry, surpassing the Standard Quantum Limit Wineland et al. (1992); Bollinger et al. (1996). The results for GHZ state preparation with this Hamiltonian are shown in Figure 6. Interestingly one of the emergent extra control Hamiltonians in Eq. 12 is $S_{y}^{2}$, which is used in the two-axis twisting Hamiltonian for the preparation of even better spin-squeezed states.
@@
As in both the short- and long-range spin models, we observe that we can prepare states with much better fidelity using these more efficient paths, though the improvement is less dramatic. The performance could be enhanced further by considering finite energy norms for the variational optimization of the action, as we briefly discuss next and as was done in a different classical model Gjonbalaj, Campbell, and Polkovnikov (2022).
IV.2 Ground State Optimization
As previously mentioned, we can replace $\operatorname{Tr}\left[G_{\lambda}^{2}\right]$ in Eq. (4) by $\operatorname{Tr}\left[G_{\lambda}^{2}\exp\left(-\beta H\right)\right]$, where $\beta$ is the inverse temperature. The original action is thus equivalent to an infinite temperature ($\beta=0$) optimization. By changing the temperature, higher weight will be assigned to lower-energy eigenstates in the optimization. As we are concerned in this paper with quantum annealing problems involving the ground state, the natural question to ask is what happens when we use a zero temperature ($\beta=\infty$) optimization, where we only optimize over transitions into and out of the ground state. Because the effect of ground state vs. infinite temperature optimization is separate from finding extra controls, in this section we only focus on the local CD driving along the original adiabatic path. One can of course combine the two approaches together, but such full optimization is outside the scope of this work. We note on passing that in some situations like in classical systems one has to deal with finite temperature actions to get meaningful results for the AGP Gjonbalaj, Campbell, and Polkovnikov (2022).
To study the effect of the ground state action, we consider the long-range model with $\alpha=2$ defined in Eq. (13). We note that if we use the short range model corresponding to $\alpha=\infty$ instead then there is no effect of temperature on the action. This fact follows from the equivalence of the short-range model to a set of independent two-level systems, where all operator norms are identical for the excited and ground states Kolodrubetz et al. (2017). In Figure 7 we plot the fidelities for the local CD protocols obtained using infinite and zero temperature actions. We see that at sufficiently high orders of the variational ansatz where the infinite temperature protocol already performs very well, the ground state optimization provides even further improvement. However, at small $\ell<3$ the ground state optimization tends to do slightly worse. A possible physical explanation is that the ground state protocol does poorly at suppressing transitions from the excited states, so if the system gets excited the infinite temperature protocol does a better job of suppressing further excitations.
Figure 7: For the annealing problem of Eq. (13) with $N=8$, we show the difference in final state fidelity between the infinite temperature (all eigenstates) and zero temperature (ground state only) optimization of the AGP. Here, $\ell$ is as defined in Eq. (5). As the order increases and the evolved state is increasingly constrained to the ground state manifold, the zero temperature optimization becomes superior.
In principle, unlike the $T=\infty$ optimization, the $T=0$ optimization requires a priori knowledge of the ground state, which would generally require a separate quantum simulator. However, we can use instead the approximate ground state prepared using the infinite-temperature CD protocol to further optimize the local AGP, and then use that new local AGP to prepare a better ground state. This procedure can be iterated, returning the approximate ground state AGP. We note that this iterative scheme does not require any extra computational resources and can be done using the same simulator (experimental or computational) used to simulate time evolution of the state.
Concretely, to implement this iterative procedure we first evolve the initial state according to the $T=\infty$ protocol. Then we use this approximate ground state along the protocol to recompute the action and find the new variational local AGP. Alternatively, one can perform experimental measurements of the operators entering the action $S_{\lambda}$ and then minimize it with respect to the variational parameters. We note that this local quadratic minimization can be always done on a classical computer. Then, the same initial state can be evolved according to this new protocol, which will produce a new wavefunction leading to the new action and new variational AGP. This procedure can be iterated, and provided that the orthonormal Krylov space construction of the AGP is used, the iterative scheme generally converges. We illustrate the results of this approximate ground state optimization in Figure 8. While the converged fidelity is slightly less than that obtained using the true ground state optimization, it still shows significant improvement if large enough values of $\ell$ are used.
Figure 8: The fidelity of the final state after annealing using the iterative optimization procedure, without any extra controls. At iteration zero, the evolution happens according to the infinite temperature protocol. Then, this is used to construct an “ground state” protocol, where the ground state is the state obtained from the previous evolution. This procedure is iterated until convergence. The dashed lines show the final state fidelity when the true ground state, obtained by exact diagonalization, is used.
In this section, we have applied this iterative procedure to local CD driving without any extra controls, i.e. with a fixed ground state path. This can also be combined with the with the extra controls protocol that we describe in this work for further improvement. This motivates the development of a joint variational principle for both the variational AGP coefficients $\alpha_{k}$ and the extra control couplings $\beta^{(n)}$. We leave the development of this principle to future work.
@@ V Conclusions & Outlook
We have proposed a systematic method for finding extra control terms for quantum annealing protocols with which to augment a Hamiltonian, so that local CD driving prepares a target ground state with exponentially better fidelity. Phrased another way, our method specifies how to find the paths through the space of possible Hamiltonian couplings along which approximate CD driving will be more efficient. We find these paths by restricting the additional terms to those which preserve the commutator structure of the original CD driving term. These terms can be effectively engineered via Floquet protocols without needing to couple directly to any new terms. We have tested this method on 1D spin chains with short- and long-range interactions and showed that it allows for a large increase in fidelity when preparing GHZ states via quantum annealing.
While this work has been principally concerned with proposing this technique and testing it with simple models, there are many others to which this recipe for more efficient paths might be applied. Going beyond one-dimensional systems, exploring models with frustration, or mappings to e.g. satisfiability problems on arbitrary graphs are all areas where this recipe may yield improvement. Another interesting direction is to pursue a joint variational principle for both the variational AGP coefficients and the extra control couplings. There are also more fundamental questions yet to be answered, such as rigorously connecting this approach and similar types of shortcuts Ibáñez et al. (2012), or similar methods in quantum dynamics such as the flow equation approach Kehrein (2006).
@@ Acknowledgments
This work was supported by NSF Grant DMR-2103658 and the AFOSR Grant FA9550-21-1-0342. The authors thank Aashish Clerk, Andrew Daley, Callum Duncan, Michael Flynn, Stefan Kehrein and Tatsuhiko Ikeda for helpful discussions. Additionally, the anonymous referees posed several interesting questions leading to further development of this work. The authors acknowledge that the computational work reported in this paper was performed on the Shared Computing Cluster administered by Boston University’s Research Computing Services. Many of the numerical computations were performed using QuSpin Weinberg and Bukov (2017, 2019). The exact code used is available online  [^1] See https://github.com/smorawetz/CD-extra-controls.git.
@@ Appendix A Construction of AGP in Krylov space
Let us provide more detail about how the approximate AGP is constructed. In particular, we use a Krylov space construction of the AGP which is slightly different to that described in Ref. Claeys et al., 2019. The main difference is that we demand that subsequent terms in the commutator expansion be orthogonal to each other. In particular, we write the AGP in terms of Krylov space operators $O_{k}$ with coefficients $\gamma_{k}$ (see also Refs. Bhattacharjee, 2023; Takahashi and del Campo, 2023)
@@
where $D=2^{N}$ is the dimension of the Hilbert space, and $H$ is the Hamiltonian. It is critical that in order to perform the ground state optimization discussed in IV.2 correctly, one must replace the trace inner product by the ground state average: $(A|B)=\langle\psi_{0}|A^{\dagger}B|\psi_{0}\rangle$. We reiterate that the iterative procedure described in the main text can obtain these without the use of a quantum computer. We construct the Krylov space operators $O_{k}$ according to the following algorithm:
@@
$S_{\lambda}=1+2\gamma_{1}b_{0}b_{1}+\sum_{k=1}^{\ell}(\gamma_{k}(b_{2k}^{2}+b_{2k-1}^{2})+2b_{2k}b_{2k+1}\gamma_{k}\gamma_{k+1})$
@@
00 0 $=\gamma_{k}A_{k}+\gamma_{k+1}B_{k}+\gamma_{k-1}B_{k-1}+b_{0}b_{1}\delta_{k,1}$
$A_{k}$ $=b_{2k-1}^{2}+b_{2k}^{2}$
@@
Appendix B Detailed Construction of AGP with Extra Controls using Floquet Pulses
[…]
Figure 9: A cartoon illustration of the kinds of a pulse sequence required for implementation of the augmented protocol. This echo-type sequence is designed in a way that different pulses almost cancel each other such that the terms appearing in different order of the Magnus expansion are of the same order.
As an illustration of this method, we show here how one can engineer the pulse sequence to generate the Floquet Hamiltonian containing the desired commutators, which appear at second order in the Magnus expansion (see below). It is conceptually straightforward to generalize this procedure to higher orders or protocols with more pulses. The Magnus expansion is controlled by the Floquet period $T$, or equivalently the Trotterization time step, which should be sufficiently short. The other time scale if the protocol time $\tau\gg T$ controls the rate of change of the coupling $\dot{\lambda}$.
We choose the strength of the pulses so that we get the terms we want in the Magnus expansion at first and second order, and as long as the expansion is written in terms of these small parameters the higher order terms in the expansion may be neglected. Then the main goal of the protocol design is to choose the pulse strengths such that the low order terms match those in the desired CD Hamiltonian. As lower order commutator terms are parametrically larger than higher order terms in the driving period $T$. In order to make them of the same order, one needs to design echo-type pulses as is routinely done in NMR literature.
Within each period of the Floquet drive we have a sequence of alternating pulses of $H_{0}$ and $H_{1}$, where the $i$-th pulse has strength $\epsilon_{i}$. We want all of the pulses to be the same size, i.e. $\epsilon_{i}=\epsilon\eta_{i}$ where $\eta_{i}\sim O(1)$ as $T\to 0$. Since we are targeting the Floquet Hamiltonian $H_{F}$ to second order with coefficients proportional to $T$, and the coefficients of second order terms in the Magnus expansion have three powers of the $\epsilon_{i}$, we take $\epsilon=T^{1/3}$.
Here we consider here a particular six-pulse sequence, with terms $\{\epsilon_{1}H_{0},\epsilon_{2}H_{1},\epsilon_{3}H_{0},\epsilon_{4}H_{1},\epsilon_{5}H_{0},\epsilon_{6}H_{1}\}$. Computing the Magnus expansion Bukov, D’Alessio, and Polkovnikov (2015) up to $O(T)$ gives:
$H_{F}T$ $=f_{0}H_{0}+f_{1}H_{1}-if_{10}[H_{1},H_{0}]$
$+f_{010}[H_{0},[H_{1},H_{0}]]+f_{110}[H_{1},[H_{1},H_{0}]]$
The coefficients $f_{i}(\lambda)$ are determined by Magnus expansion, and have been worked out in full for an arbitrary number of pulses and higher order terms  [^2] For code to calculate these, use Mathematica notebook at https://github.com/smorawetz/CD-Floquet-pulses.git. The first few terms are
$f_{0}$ $=\epsilon_{1}+\epsilon_{3}+\epsilon_{5},\quad f_{1}=\epsilon_{2}+\epsilon_{4}+\epsilon_{6}$
$f_{01}$ $=\epsilon_{2}{\epsilon_{1}+\epsilon_{3}+\epsilon_{5}\over 2}+\epsilon_{4}{\epsilon_{1}+\epsilon_{3}-\epsilon_{5}\over 2}+\epsilon_{6}{\epsilon_{1}-\epsilon_{3}-\epsilon_{5}\over 2}$
$\vdots$
We will now match this expansion with $\tilde{H}_{\rm CD}T$. We assume that $\lambda$ changes sufficiently slowly so that it is nearly constant within a Floquet period. The CD Hamiltonian that want to match is:
$\tilde{H}_{\rm CD}=\lambda H_{0}+(1-\lambda)H_{1}-i\dot{\lambda}\alpha_{1}[H_{1},H_{0}]$
$+\beta^{(1)}[H_{0},[H_{1},H_{0}]]+\beta^{(2)}[H_{1},[H_{1},H_{0}]]$
so that the corresponding terms in the Magnus expansion are
$f_{0}=\lambda,f_{1}=(1-\lambda),f_{01}=\dot{\lambda}\alpha_{1}(\lambda)$
$f_{010}=\beta^{(1)}(\lambda),f_{110}=\beta^{(2)}(\lambda)$
Writing the pulse strengths as $\epsilon_{i}=T^{1/3}\eta_{i}$, the following protocol can be used:
$\eta_{1}$ $=-\frac{\beta_{1}^{2/3}}{\beta_{2}^{1/3}}+\sqrt{\frac{\alpha\beta_{1}}{3\beta_{2}}}\dot{\lambda}^{1/2}T^{1/6}+{1\over 3}(1-\lambda)T^{2/3}$
$\eta_{2}$ $=\frac{\beta_{2}^{2/3}}{\beta_{1}^{1/3}}+\sqrt{\frac{3\alpha\beta_{2}}{\beta_{1}}}\dot{\lambda}^{1/2}T^{1/6}+{1\over 3}\lambda T^{2/3}$
$\eta_{3}$ $=2\frac{\beta_{1}^{2/3}}{\beta_{2}^{1/3}}+\sqrt{\frac{\alpha\beta_{1}}{3\beta_{2}}}\dot{\lambda}^{1/2}T^{1/6}+{1\over 3}(1-\lambda)T^{2/3}$
$\eta_{4}$ $=\frac{\beta_{2}^{2/3}}{\beta_{1}^{1/3}}+{1\over 3}\lambda T^{2/3}$
$\eta_{5}$ $=-\frac{\beta_{1}^{2/3}}{\beta_{2}^{1/3}}-2\sqrt{\frac{\alpha\beta_{1}}{3\beta_{2}}}\dot{\lambda}^{1/2}T^{1/6}+{1\over 3}(1-\lambda)T$
$\eta_{6}$ $=-2\frac{\beta_{2}^{2/3}}{\beta_{1}^{1/3}}-\sqrt{\frac{3\alpha\beta_{2}}{\beta_{1}}}\dot{\lambda}^{1/2}T^{1/6}+{1\over 3}\lambda T^{2/3},$
This choice of the protocol guarantees that $H_{F}$ and $\tilde{H}_{\rm CD}$ coincide up to the terms scaling as higher powers of $T$, which can be made arbitrary small by taking the limit $T\to 0$.
Appendix C Performance of YY vs. ZXZ controls in Short-Range Model
@@ Appendix B Detailed Construction of AGP with Extra Controls using Floquet Pulses
In Figure 10, we show a contour plot of the final fidelity for different combinations of $\beta_{YY}$ and $\beta_{ZXZ}$ using local CD driving with $\ell=1$ and $\ell=2$ as defined in Eq. (5) for the Hamiltonian above with $N=10$. We find that for $\ell=1$ the optimal protocol is very close to one where only one of the controls is used, i.e. where either $\beta_{YY}=0$ or $\beta_{ZXZ}=0$. The situation reverses for $\ell=2$, where the optimal performance is achieved for $\beta_{ZXZ}\approx-\beta_{YY}$. Figure 3 in the main text shows the protocol performance along the corresponding optimal directions.
Figure 10: The fidelity obtained by augmenting the short-range GHZ state preparation for $N=10$ by either $YY$ or $ZXZ$ controls. The origin represents the fidelity of state preparation without any extra controls. The top shows $\ell=1$ an the bottom shows $\ell=2$. The x and y axis show the coefficient for the $YY$ and $ZXZ$ controls respectively. From the symmetry of this plot, it is clear that neither the $YY$ or $ZXZ$ term has an advantage over the other. Starting at $\ell=2$, using both simultaneously provides a significant advantage.
Appendix D Detailed Steps to Implement the Method
In this section, we will describe in detail the step-by-step process by which we find these improved paths for local CD driving (see also Ref. Čepaitė et al., 2023). The first step is to determine the Hamiltonian for the physical system of interest, and represent it as a matrix. Furthermore, it is useful to leverage any symmetries present in the system and so construct the Hamiltonian in the relevant symmetry sector, which can be done easily with QuSpin or equivalent software package. This will make the local CD driving more efficient by reducing the number of transitions it tries to suppress. In the language of Eq. (6), this means determining in $H_{0}$ or $H_{1}$ and then forming
@@ Appendix D Detailed Steps to Implement the Method
We then compute the fidelity using Eq. (10). We employ the Powell minimization algorithm Powell (1964) to choose the values of $\beta^{(1)}$ and $\beta^{(2)}$ which maximize the fidelity. Practically, we limit the values the coefficients can take to $|\beta^{(i)}|<3$. This concludes the steps to implementing our protocol. In Figure 11 we apply this procedure to the short-range model Hamiltonian of Eq. (11). This is the same as Figure 3 except using the $H_{c}^{(1)}$ and $H_{c}^{(2)}$ of Eq. 12 instead of $YY$ and $ZXZ$.
Figure 11: Improvement in the final state fidelity obtained by using the extra control ansatz $H_{c}^{(1)}=[H_{0},[H_{1},H_{0}]]$ and $H_{c}^{(2)}=[H_{1},[H_{1},H_{0}]]$ to prepare a GHZ state by annealing the short-range Hamiltonian of Eq. (11). This can be compared with Figure 3 where we use $H_{c}^{(1)}=YY$ and $H_{c}^{(2)}=ZXZ$ for the same protocol. Note that they are extremely similar because the corresponding extra controls are linear combinations of each other up to a rescaling of $ZZ$ and $X$ in the original Hamiltonian. As before, red points indicate following the “naive” original path, whereas blue indicates following an “augmented” path.
@@
Appendix E Explanation of Different Regimes of Fidelity Improvement
Figure 12: A snapshot of the variational optimization in terms of fitting excitation frequencies, for the short-range model with $N=6$. On top is the “naive path” of the original Hamiltonian, and on the bottom is the augmented path. Although this is just one snapshot, it is apparent from this that the extra control Hamiltonian $H_{c}=YY$ makes local CD driving more efficient by bringing multiple excitation frequencies together.
As remarked upon in the main text, in Figures 3, 5 and 11, we see two types of behavior for the fidelity of the final state prepared with extra control Hamiltonians. The first regime is a plateau where we can prepare the final state with near unit fidelity up to a certain system size, whereas the second regime has less than unit fidelity but is still exponentially improved when compared to the “naive” path.